OpenStudy (anonymous):

The measure of angle A is twice that of angle B and one less than that of angle C. If the sum of the measures of the three angles is at least 66, what is the least possible measure of angle A?

OpenStudy (anonymous):

Need the solution. :)

OpenStudy (anonymous):

Wanna learn how to do it or just the answer?

OpenStudy (anonymous):

i wanna learn how to do it. :) please?

OpenStudy (anonymous):

Sure, you'll need a pen and a paper to make it simpler for you tho :)

OpenStudy (anonymous):

okay. :) then?

OpenStudy (anonymous):

First off.. all those kinds of mathematical problems have one huuuuge key to them : you have to be able to translate those words into mathematical sentences (ie: with numbers)

OpenStudy (anonymous):

1st, draw a square... and name 3 corners : A, B and C... it'll help visualize the problem beofre we attempt it

OpenStudy (anonymous):

Basically we want to know how to represent the values of A, B and C... once that's done, the rest unfolds

OpenStudy (anonymous):

To start, we'll name the first corner A, and make it the value X and go from there...

OpenStudy (anonymous):

so... A = x

OpenStudy (anonymous):

Can you tell me what we know about B?

OpenStudy (anonymous):

A = 2B A = C-1 2B = C - 1 C-2B = 1 A + B + C is at least 66?

OpenStudy (anonymous):

how will i know the value of a b and c?

OpenStudy (anonymous):

still there?

OpenStudy (anonymous):

yes, still here, i think my internet froze for a sec

OpenStudy (anonymous):

Hmm... Can you explain to me your 2B = C-1?

OpenStudy (anonymous):

i think the value of a is c-1

OpenStudy (anonymous):

Well I'm realizing I have one way to solve this, using an equiation with "x" as a variable... But is that the way you,re supposed to use too?

OpenStudy (anonymous):

Like... I'd made an equation using the \[\ge\] symbol

OpenStudy (anonymous):

Like... I'd try to describe all my angles... I know that A is twice the value of B... so... if we say B = x degrees then A = 2x degrees makes sense?

OpenStudy (anonymous):

Then you do on the square you drew you write "x" inside the corner of B to mark it as it's value... and write "2x" in A's corner to mark it's value

OpenStudy (anonymous):

okay. got it. then? :)

OpenStudy (anonymous):

So far we have... B = x A = 2x C = ? Let's find C! What do we know of C? ... well we know that the measure of A is one less than that of angle C... How would you write that mathematically?

OpenStudy (anonymous):

Do you know what I mean by "write that mathematically"?

OpenStudy (anonymous):

A =C-1 right?

OpenStudy (anonymous):

or C=2x-1

OpenStudy (anonymous):

Hmm, incorrect... read the sentence again ;) gotta be careful! "A is one less than C"...

OpenStudy (anonymous):

C=1-2x?

OpenStudy (anonymous):

hello? :)

OpenStudy (anonymous):

It's kind of counter intuitive... that's why you have to be careful... If they say A is one less than C... A = 2x <--- this is LESS than C... so C is one more than A... how do you write "one more" in math? ... +1! C is actually A+1

OpenStudy (anonymous):

"A is one less than C" can also be read "C is one more than A". Do you agree that it means the same thing? It makes the understanding simpler... And avoids small mistakes like that

OpenStudy (anonymous):

Makes sense?

OpenStudy (anonymous):

okaaay. So B=x A=2x and C= 2x+1 :)

OpenStudy (anonymous):

yap.

OpenStudy (anonymous):

Yep yep! Now that we have this... rest is just putting it together... Whats the last part of the question? How do you put "the sum of the measures of the three angles is at least 66" in a math sentence?

OpenStudy (anonymous):

all right then? :)

OpenStudy (anonymous):

Well tell me? How do you write "the sum of the measures of the three angles is at least 66"? This will be our equation to solve

OpenStudy (anonymous):

oooh. thanks for that! :) all's clear.

OpenStudy (anonymous):

Cool... so what is the answer?

OpenStudy (anonymous):

\[x+2x+2x+1\ge66?\]

OpenStudy (anonymous):

hot stuff :) so what is x?

OpenStudy (anonymous):

oooh. thanks for that! :) all's clear.

OpenStudy (anonymous):

13?

OpenStudy (anonymous):

oooh. thanks for that! :) all's clear.

OpenStudy (anonymous):

is it right?

OpenStudy (anonymous):

Good, although an exact answer would be \[x \ge 13\] Sometimes teachers can be picky with stuff like that... :P

OpenStudy (anonymous):

you've solved the equation, but not the problem yet... What is the question asked in the problem?

OpenStudy (anonymous):

oooh. thanks for that! :) all's clear.

OpenStudy (anonymous):

thaaanks! Got that already. I just need to substitute it right?

OpenStudy (anonymous):

Yep... Just be careful cuz when i solved it quickly at 1st I was gonna answer "at least 13", but that wouldn,t be correct... since we have to substitute in A, and A=2x... A= 2x and x larger or equal as 13 so \[A \ge2*13\] \[A \ge26\]

OpenStudy (anonymous):

oooh. thanks for that! :) all's clear.

OpenStudy (anonymous):

so the answers are 13, 26, and 17?

OpenStudy (anonymous):

Well those are the minimal values yes... but the actual question is "what is the least possible measure of angle A?"

OpenStudy (anonymous):

to which the answer is 26 ;)

OpenStudy (anonymous):

Good work! hihihi

OpenStudy (anonymous):

Best advice I could give you is to split the question is different pieces and make sure you understand it 100%, you seem to read it too quickly and might lose points in an exam for answering the wrong question or something ;)

OpenStudy (anonymous):

oooh. thanks for that! :) all's clear.

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

My pleasure! Alright, take care! :)